What is the solution to 5x - 1/2x = 18 please show the work.
What is the solution to -8x -2x = 15, please show the work.

Answer by checkley71(8403)

5X-1/2X=18
4.5X=18
X=4
PROOF
5*4-1/2*4=18
20-2=18
18=18
---------------------------------
-8X-2X=15
-10X=15
X=15/-10
X=-1.5
PROOF
(-8*-1.5)-(2*-1.5)=15
12+3=15
15=15